# Pretraining Scaling Laws for Generative Evaluations of Language Models

### v1

@@ Appendix B Methodology: Number of Samples Per Problem Per Model
Because sampling from language models is expensive and because our compute budget is limited, we needed to judiciously choose how many samples to draw per problem per model. A key consideration is that the number of samples per problem sets the resolution, i.e., if we draw $n$ samples per problem, then any pass rate on that problem below $1/n$ will likely appear to be 0. Since pass rates typically increase with pretraining compute, we chose a loose heuristic based on the pretraining compute of each model. For a model with $C$ pretraining FLOP, we aimed to sample
$\text{Num. Samples per Problem }=\operatorname{clip}(100\cdot 10^{23-\log_{10}(C)},3.2e4,1e6).$
$3.2e4$ was chosen so that we would have more samples per problem than the largest number of attempts per problem $k$ we considered $(1e4)$. Fig. 8 shows how many samples per problem per model we had at the time our analyses were conducted. Additional sampling would have likely benefited our analyses.
Figure 8: Number of Samples per Model per Problem. This heatmap visualizes how many samples we had drawn per model (y-axis) per problem (x-axis) at the time our analyses were conducted.
Appendix C Scaling Law Parameters by Number of Attempts Per Problem $k$
For the three scaling laws we consider, we visualize how the fit parameters change as a function of the number of attempts per problem $k$ (Fig. 9, Fig. 10, Fig. 11).
Figure 9: Pretraining Compute Scaling Law Parameters by Number of Attempts per Problem (Full Fits). We visualize how the parameters of Eqn. 3 change with respect to $k$. Hue corresponds to $k$.
Figure 10: Model Parameters and Tokens Scaling Law Parameters by Number of Attempts per Problem (Full Fits). We visualize how the parameters of Eqn. 4 change with respect to $k$. Hue corresponds to $k$.
Figure 11: Gold Reference Log Likelihoods Scaling Law Parameters by Number of Attempts per Problem (Full Fits). We visualize how the parameters of Eqn. 5 change with respect to $k$. Hue corresponds to $k$.
Appendix D Relation Between the Compute Scaling Law and the Parameters + Tokens Scaling Law
@@
Form the Lagrangian $\mathcal{L}=AN^{-\beta}+BD^{-\gamma}+\lambda\,(ND-C/c)$. First-order conditions give $\beta AN^{-\beta}=\gamma BD^{-\gamma}$ and $ND=C/c$, from which (A.5) follows. Substituting $N^{*},D^{*}$ into $F(N,D)$ yields $\mathcal{F}(C)=C_{0}(k)\,C^{-\alpha(k)}$ with $\alpha(k)$ and $C_{0}(k)$ as in Eqs. 11-12, and $E(k)=\mathcal{E}_{0}(k)$.
@@
    All quantities may depend on $k$; the mapping in Eqs. 11-13 applies point-wise in $k$. The conversion constant $c$ enters only via $C_{0}$, not $\alpha$.
D.1 Multiplicative deviation from the optimal compute envelope

### v2

@@ Appendix B Methodology: Number of Samples Per Problem Per Model
Because sampling from language models is expensive and because our compute budget is limited, we chose a compute-aware sampling strategy designed to balance resolution with cost. A key consideration is that the number of samples per problem sets the resolution, i.e., if we draw $n$ samples per problem, then any pass rate on that problem below $1/n$ will likely appear to be 0.
For each model checkpoint, we took the following strategy: we drew a minimum of $2^{14}$ samples per model per problem, and then continued sampling until $10$ successes were obtained or until a maximum of $2^{15}$ samples were drawn. $2^{14}$ was chosen so that we would have more samples per problem than the largest number of attempts per problem $k$ we considered $(1e4)$.
Fig. 9 shows how many samples per GSM8K problem per model we had at the time our analyses were conducted, and Fig. 10 shows how many samples per MATH problem per model we had at the time our analyses were conducted. In total, we had $\sim 500M$ total samples for GSM8K and $\sim 400M$ total samples for MATH. Additional sampling would have likely benefited our analyses.
Figure 9: Number of Samples per Model per GSM8K Problem. This heatmap visualizes the number of samples per model (y-axis) per problem (x-axis) at the time our analyses were conducted.
Figure 10: Number of Samples per Model per MATH Problem. This heatmap visualizes the number of samples per model (y-axis) per problem (x-axis) at the time our analyses were conducted.
Appendix C Estimation of Total Sampling Cost
Sampling was performed using vLLM (Kwon et al., 2023). We generated 128 samples of the same problem at a time to make use of prefix caching for more efficient sampling.
C.1 Estimation of GSM8K Sampling Costs
Estimates of the GPU time per 128 samples per model size for GSM8K are provided in Table 1.
Pythia Model Parameters A100-s Per 128 Samples (Approximate)
14M 4
70M 4
160M 5
410M 7
1B 9
1.4B 11
2.8B 13
6.9B 25
12B 36
Table 1: A100-Seconds Per 128 Samples by Model Size on GSM8K Problems.
At a minimum of $3.2e4$ samples per problem, $128$ problems per model checkpoint, $8$ model checkpoints per model size, our total compute costs are lower bounded by:
[…]
$\geq 292e6\text{ A100-s}$
$\geq 81e3\text{ A100-hr}$
Thus, sampling our GSM8K data is lower bounded by $81e3$ A100-hr. To estimate the total financial cost of this sampling, we looked up pricing of compute providers. As of November 2025, Lambda offers A100s at $\mathdollar 1.79$/A100-hr, RunPod at $\mathdollar 1.39$/A100-hr, Tensordock at $\mathdollar 1.63$/A100-hr, and Vast-AI at $\mathdollar 1.40$/A100-hr. In total, the financial cost of our GSM8K sampling data is more than $\mathdollar 112k$.
C.2 Estimation of MATH Sampling Cost
Estimates of the GPU time per 128 samples per model size for MATH are provided in Table 2.
Pythia Model Parameters A100-s Per 128 Samples (Approximate)
14M 4
70M 4
160M 5
410M 8
1B 10
1.4B 10
2.8B 15
6.9B 27
12B 39
Table 2: A100-Seconds Per 128 Samples by Model Size on MATH Problems.
Appendix D GSM8K Scaling Law Parameters by Number of Attempts Per Problem $k$
For the three scaling laws we consider, we visualize how the fit GSM8K scaling law parameters change as a function of the number of attempts per problem $k$ (Fig. 11, Fig. 12, Fig. 13).
Figure 11: GSM8K Pretraining Compute Scaling Law Parameters by Number of Attempts per Problem (Full Fits). Scaling law parameters of Eqn. 3 as functions of $k$. Hue corresponds to $k$.
Figure 12: GSM8K Parameters and Tokens Scaling Law Parameters by Number of Attempts per Problem (Full Fits). Scaling law parameters of Eqn. 4 as functions of $k$. Hue corresponds to $k$.
Figure 13: GSM8K Gold Reference Log Likelihoods Scaling Law Parameters by Number of Attempts per Problem (Full Fits). Scaling law parameters of Eqn. 5 as functions of $k$. Hue corresponds to $k$.
Appendix E Relation Between the Compute Scaling Law and the Parameters + Tokens Scaling Law
@@
Form the Lagrangian $\mathcal{L}=AN^{-\beta}+BD^{-\gamma}+\lambda\,(ND-C/c)$. First-order conditions give $\beta AN^{-\beta}=\gamma BD^{-\gamma}$ and $ND=C/c$, from which (A.5) follows. Substituting $N^{*},D^{*}$ into $F(N,D)$ yields $\mathcal{F}(C)=C_{0}(k)\,C^{-\alpha(k)}$ with $\alpha(k)$ and $C_{0}(k)$ as in Eqs. 12-13, and $E(k)=\mathcal{E}_{0}(k)$.
@@
    All quantities may depend on $k$; the mapping in Eqs. 12-14 applies point-wise in $k$. The conversion constant $c$ enters only via $C_{0}$, not $\alpha$.
E.1 Multiplicative deviation from the optimal compute envelope
@@
Appendix F How Scaling Depends on the Benchmark: Three Pretraining Scaling Laws for MATH
In the main text, we presented results for GSM8K (Cobbe et al., 2021). Here, we present corresponding results for MATH (Hendrycks et al., 2021). For a focused discussion of the similarities and differences between the two benchmarks, please see Section 7 in the main text.
F.1 Fitting and Predicting Pass Rates from Pretraining Compute
Figure 14: Scaling of MATH Pass Rates with Pretraining Compute (Full Fit). Each panel fits Eqn. 3 $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)(C,k)=E_{0}(k)+C_{0}(k)\cdot C^{-\alpha(k)}$ to MATH pass rates for Pythia checkpoints across $\sim$5 orders of magnitude of pretraining compute for $k\in\{1,10^{2},10^{4}\}$.
Figure 15: Predicting MATH Pass Rates from Scaling Pretraining Compute (Backtesting). We evaluate predictability via backtesting: we iteratively fit Eq. 3 on subsets of models up to a maximum pretraining compute, then extrapolate to predict the most expensive model’s $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)$. The most expensive model is Pythia 12B-parameter 300B-token ($\approx 2.16\times 10^{22}$ FLOP).
F.2 Fitting and Predicting Pass Rates from Parameters and Tokens
Figure 16: Scaling of MATH Pass Rates with Parameters and Tokens (Full Fit). Each panel fits Eqn. 4, $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)(N,D,k)=\mathcal{E}_{0}(k)+N_{0}(k)\,N^{-\beta(k)}+D_{0}(k)\,D^{-\gamma(k)}$.
Figure 17: Predicting MATH Pass Rates from Scaling Parameters and Tokens (Backtesting). We evaluated how accurately the parameters + tokens scaling law (Eqn. 4) predicts the most expensive model’s $\log(\mathrm{pass}_{\mathcal{B}}@k)$.
F.3 Fitting and Predicting Pass Rates from Gold Reference Likelihoods
Figure 18: Scaling of MATH Pass Rates with Gold Reference Likelihood (Full Fit). Each panel fits Eqn. 6, $\log(\mathrm{pass}_{\mathcal{B}}@k)=\xi_{0}+K_{0}(k)\cdot[-\log\!\big(\mathrm{GoldProb}_{\mathcal{B}}\big)]^{\kappa(k)}$, for $k\in\{1,10^{2},10^{4}\}$.
Figure 19: Predicting MATH Pass Rates from Gold Reference Likelihoods (Backtesting). We evaluated how accurately the gold reference likelihoods scaling law (Eqn. 6) predicts the most expensive model’s $\log(\mathrm{pass}_{\mathcal{B}}@k)$.
Appendix G MATH Scaling Law Parameters by Number of Attempts Per Problem $k$
For the three scaling laws we consider, we visualize how the fit MATH scaling law parameters change as a function of the number of attempts per problem $k$ (Fig. 20, Fig. 21, Fig. 22).
Figure 20: MATH Pretraining Compute Scaling Law Parameters by Number of Attempts per Problem (Full Fits). Scaling law parameters of Eqn. 3 as functions of $k$. Hue corresponds to $k$.
Figure 21: MATH Parameters and Tokens Scaling Law Parameters by Number of Attempts per Problem (Full Fits). Scaling law parameters of Eqn. 4 as functions of $k$. Hue corresponds to $k$.
Figure 22: MATH Gold Reference Log Likelihoods Scaling Law Parameters by Number of Attempts per Problem (Full Fits). Scaling law parameters of Eqn. 5 as functions of $k$. Hue corresponds to $k$.